DogGirl27:

What is the first step in solving the quadratic equation -5x2+8=133? taking the square root of both sides of the equation subtracting 8 from both sides of the equation squaring both sides of the equation adding 8 to both sides of the equation

GoodFellow:

You need to isolate the squared term on one side, therefore, you need to do the second one, first

Alex1513:

The first step is subtract 8 from both sides of the equation Then you take the square root of both sides of the equation Finally you divide from both sides op the equation by -5 -5x2+8=133 -5x2 + 8 - 8 = 133 - 8 -5x2 = 125 -5x = √125 -5x = 11.18 -5x / -5 = 11.18/-5 x = 2.236

Alex1513:

subtracting 8 from both sides of the equation

DogGirl27:

thanks so much

Alex1513:

I SHOWED WORK

DogGirl27:

extrinix do you have to be such a jerk

DogGirl27:

if you dont like it then leave

DogGirl27:

well you dont need to parent us thanks boo

DogGirl27:

if your going to correct me then go help someone else bc i dont need that